OpenStudy (anonymous):

HELP Find the solution to the system algebraically. Show all work for full credit. Leave your answers in fraction form. y=6x^2 7x+y=3

OpenStudy (whpalmer4):

Solve second equation for y in terms of x. Replace y in first equation with that, giving you a quadratic, and solve by your favorite method.

OpenStudy (anonymous):

so how would you set this up as

OpenStudy (whpalmer4):

7x+y=3 y=3-7x 6x^2=3-7x 6x^2+7x-3=0 Solve...

OpenStudy (anonymous):

so basically u plug y=6x^2 to y on the equation below which is 7x+6x^2=3 adn then you organize it so its 6x^2+7x=3 and then u move the 3 to other side which becomes -3 so

OpenStudy (anonymous):

basically u gotta facot 6x^2+7x-3 i would like to give you an answer but i wish you to try it out before i give you the answer

OpenStudy (anonymous):

is it 40 ?

OpenStudy (whpalmer4):

Don't give them the answer, help them work it out. Much more helpful in the long run.

OpenStudy (anonymous):

he didn t give me the answer ??

OpenStudy (whpalmer4):

No, a quadratic has two unique solutions unless it is a perfect square. This is not, and 40 is incorrect. Do you get a true number statement if you square 40, multiply by 6, add 7*40 to that and subtract 3? If that equals 0, you have a solution. Or a really big value of 0 :-)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I got 9877

OpenStudy (whpalmer4):

He said he wanted to give you the answer. Do you know the quadratic formula solution? You have a = 6, b = 7, c =-3 \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (whpalmer4):

That you got 9877 means 40 isn't the solution!

OpenStudy (whpalmer4):

If 40 was the solution, you would get 0. We are trying to find values of x such that \[6x^2+7x-3=0\]

OpenStudy (anonymous):

how would you find the values if x

OpenStudy (anonymous):

do you solve for x

OpenStudy (whpalmer4):

Use the formula I gave you with the a,b,c

OpenStudy (anonymous):

.024316339???

OpenStudy (anonymous):

-7+7^2-(6)(-3)/(2)6

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

@xoxox123 : Have a look @ attachment bro U had tried your best

OpenStudy (anonymous):

thats not the same problem.. @Anu2401

OpenStudy (anonymous):

Thats same brother

OpenStudy (whpalmer4):

@xoxox123 you forgot to do the square root and a multiplication when you did the formula, \[x=\frac{-7\pm\sqrt{7^2-4(6)(-3)}}{2(6)}=\frac{-7\pm\sqrt{121}}{12}=\frac{-7\pm11}{12}\]\[x=\frac{1}{3}, x=-\frac{3}{2}\]